OpenStudy (anonymous):

let U be a square matrix such that U transpose*U = I. show that |det U| =1

OpenStudy (zarkon):

\[1=Det(I)=Det(U^'U)=Det(U')Det(U)=Det(U)Det(U)=Det(U)^2\] \[Det(U)=\pm\sqrt{1}=\pm 1\] thus \[|Det(U)|=1\]

OpenStudy (anonymous):

thanks